# AlpacaFarm: A Simulation Framework for Methods that Learn from Human Feedback

### 06 Dec 2023 v3

@@ 4 Validating the AlpacaFarm simulator
Figure 3: The ranking of methods trained and evaluated in AlpacaFarm matches that of methods trained and evaluated in the human-based pipeline. Each point represents one method $M$ (e.g. PPO). The x-axis shows the simulated evaluation (win-rates measured by $p_{\text{sim}}^{\text{eval}}$) on methods trained in simulation $M_{\text{sim}}$. The y-axis shows human evaluation (win-rates measured by $p_{\text{human}}$) on methods trained with human feedback $M_{\text{human}}$.
@@ 4.1 Experimental details
As a baseline and starting point for LPF methods, we fine-tuned LLaMA 7B on the 10k SFT split. We take SFT 10k to be the starting point for all LPF methods and collect the simulated preference $p_{\text{sim}}^{\text{ann}}$ and human preference $p_{\text{human}}$ from SFT 10k’s outputs (with temp=1.0) on the 10k instruction PREF split.
Then, for each of the six reference LPF methods $M$:
@@
  • We trained a few models $M$ on human preferences across hyperparameter ranges identified in simulation, evaluating the resulting model $M_{\text{human}}$ against Davinci003 with humans $p_{\text{human}}$.
In addition to the six methods, we also evaluate existing standard instruction following and base models: GPT-4 (gpt-4-0314), ChatGPT (gpt-3.5-turbo-0301), Davinci001 (text-davinci-001), LLaMA 7B (Touvron23, ), and Alpaca 7B (Taori23, ). Alpaca 7B is a LLaMA 7B model finetuned on the concatenation of all data splits, denoted SFT 52k. For these models, we measure both the simulated win-rate $p_{\text{sim}}^{\text{eval}}$ and human win-rate $p_{\text{human}}$.
At inference time, for all systems except best-of-$n$, we sample with temp=0.7 and set the maximum number of tokens to be 300. For best-of-$n$ sampling, we found a higher temperature to be helpful in encouraging output diversity, and so we rerank samples from SFT 10k with temp=1.0. We provide more thorough experimental details and hyperparameters for all methods in Appendix B.
@@
We collected reference human annotation by showing crowdworkers two potential outputs $y_{0}$ or $y_{1}$ for a given instruction $x$ and asked them to select the index $z\in\{0,1\}$ of their preferred output. Annotators are recruited from Amazon Mechanical Turk using a qualification test of 25 questions. Out of an initial pool of 34 annotators, we selected the 16 whose agreement rate was higher than 70% with the author’s annotations. We paid the annotators a median hourly rate of \$21, leading to a one-time \$3000 cost of annotating our PREF split and a recurring \$242 cost for evaluating a single model on the 805 evaluation instructions. See Appendix D for additional details including the annotation interface we provided.
@@ 4.2 End-to-end validation of AlpacaFarm
Inspecting these results more closely, we point out the two rank mismatches. The first comparison is SFT10k against SFT52k, where human annotators preferred SFT10k (44.3% vs 40.7%) while the simulator had the opposite preference (36.7% vs 39.2%, Table 2). The other mismatch is ChatGPT against PPO, where human annotators preferred PPO (55.1% vs 52.9%) unlike the simulator (46.8% vs 61.4%). In both cases, these are not major mistakes, as we do not expect SFT52k to be much worse than SFT10k or for a 7B LLaMA model to substantially outperform ChatGPT.
@@ 4.3 Validating the pairwise preferences component
We now show that modeling annotator variability in the simulated annotator ($p_{\text{sim}}^{\text{ann}}$) is necessary to capture important qualitative features of LPF model training. To do so, we compare the behavior of three of the best-performing models trained under $p_{\text{sim}}^{\text{ann}}$ with those trained using the single GPT-4 prompt $p_{\text{sim}}^{\text{GPT-4}}$, which has higher human agreement but little annotator variability.
Figure 5(c) shows the learning dynamics of these models for pairwise feedback by $p_{\text{human}}$ (left), $p_{\text{sim}}^{\text{ann}}$ (middle), and $p_{\text{sim}}^{\text{GPT-4}}$ (right) as the PPO iteration count and rerank sample counts for best-of-$n$ and expert iteration are increased. For both the human and AlpacaFarm preferences, models that more effectively optimize the surrogate reward (x-axis) improve in win-rate (y-axis) up until a point, where overoptimization of the reward occurs and win-rates decrease. In contrast, simple GPT-4 feedback shows no overoptimization, leading to the false conclusion that the LPF methods can be optimized for much longer than is actually the case. For example, Figure 5(c) right shows best-of-$1024$ to be much better than PPO, which disagrees strongly with the results on human data.
Noted in prior work Gao22 , overoptimization is the result of the proxy reward model $\hat{R}_{\phi}$ being an imperfect estimate of the true reward $R$. We hypothesize that the strong overoptimization seen for human annotators is partly due to the (inter- and intra-) variability of their annotations, which degrades the quality of the reward model. To test this hypothesis, we measured the variance of each of the three annotators by calculating the average error of a held-out annotation to the majority vote over 3 random draws. At 0.26 for $p_{\text{sim}}^{\text{eval}}$ and 0.43 for $p_{\text{sim}}^{\text{ann}}$, AlpacaFarm is close to the human variance of 0.35; in contrast, the GPT-4 annotator has a much lower variance at 0.1. Finally, in Section E.1 we ablate the AlpacaFarm design more finely and find that the added label noise provides the majority of the benefit for inducing overoptimization. Appendix C contains further analyses of bias and variability of annotators.
@@ 4.4 Validating the evaluation protocol
Finally, we test our evaluation data that combines existing open-source evaluation datasets. While we have observed that this dataset is diverse (see Figure 2), it is unclear whether it evaluates performance for any type of real-world human usage. To resolve this, we measure method-level correlations against a set of real user interactions recorded on the Alpaca Demo (Taori23, ). We manually went through the interactions and identified 200 instructions that do not contain any personal identifying information, toxic or unsafe questions, and those that refer to the chatbot directly (e.g. “who are you developed by?”). The terms of use for the demo do not allow us to publicly release this data, but we use this data to evaluate the proposed evaluation set.
@@ 5 Benchmarking reference methods on the AlpacaFarm
Table 2: AlpacaFarm evaluation results on baseline and LHF methods. win-rates are computed against Davinci003 (which achieves 50% win-rate against itself). We omitted Quark and Binary Reward Conditioning for human evaluation because they consistently underperform SFT 10k in development.
@@
Davinci001 $24.4\pm 1.5$ $32.5\pm 1.6$
LLaMA 7B $11.3\pm 1.1$ $6.5\pm 0.9$
We now study the performance of reference methods in AlpacaFarm. Table 2 contains the details of the main evaluation results (presented graphically in Figure 3). In the rest of this section, we discuss our findings from these results, demonstrating that the conclusions we reach using human feedback could have been derived using AlpacaFarm at a substantially lower cost.
@@ 5.1 Comparing LPF methods
Among the LPF methods we study, PPO performs the best in both the simulator ($47\%$) and on human feedback data ($55\%$). Notably, with a win-rate of $55\%$, the PPO trained with human feedback was preferred to ChatGPT by our human annotators for single-turn instruction-following.
@@
Even though expert iteration trains with best-of-$n$ outputs, it trails best-of-$n$ with a 3-6% win-rate gap in both simulated and human feedback workflows. This result suggests that acquiring the improvement from best-of-$n$ sampling is not straightforward and points to why more complex learning algorithms like PPO might be useful.
For Quark, we observe that although the rewards of samples improve during training, this did not result in models that perform better than the baseline SFT 10k. See Appendix B for details of our tuning.
@@ 6 Related work
Constitutional AI Bai22 simulates human feedback with AI feedback for model development to improve harmlessness and helpfulness. AlpacaFarm, on the other hand, simulates human feedback with API LLMs so that simulated experiments reflect the outcomes of experiments performed with real human feedback. Due to the difference in goals, the construction and usage of the feedback simulator are different in the two settings. For example, AlpacaFarm’s simulator perturbs LLM preferences with bit-flip label noise to mimic the noisiness of human annotation, whereas Constitutional AI’s simulator doesn’t inject extra noise.
@@ 7 Limitations and future directions
Difference between GPT4 and human feedback.
Our sandbox assumes that LLMs can simulate human feedback. Section 4 shows that this assumption holds in the sense that LLM annotations align with the mode of human preferences and replicate its many features. However, we also observe that no single LLM-based annotator captures the heterogeneity of human annotation, and substantial amounts of noise had to be injected in the simulated preference for rankings of methods trained in AlpacaFarm to match those trained with real human feedback.
In addition, we found that suitable hyperparameters for learning algorithms can be different for training with simulated feedback compared to human feedback. For example, due to changes in the scale of values of the surrogate reward model, the range of suitable KL regularization coefficients for RLHF is different. This suggests that the AlpacaFarm simulator, as it currently is, would not always be suitable to aid hyperparameter tuning to improve performance on human evaluation.
Finally, we note that our preference simulator is validated with respect to our recruited pool of crowdworkers, and mimics their preference variability as a result. The best prompts for transferring insights to other, more consistent crowdworker pools may involve lower levels of noise or a smaller prompt ensemble.
Limitation in current experiments.
Due to resource constraints (both compute and human annotation), we have not conducted extensive hyperparameter tuning for the methods we studied in Section A. For each method, we have performed at least 3 tuning runs, and we detail our tuning process in Appendix B. It is possible that with more careful tuning, the comparison between PPO and alternative approaches like expert iteration and Quark might change.
In addition, our experiments were conducted with models and datasets of a particular scale. Qualitative findings may change as we move to larger models or datasets.
Our study focuses on comparisons of methods that learn from pairwise feedback. We do not study other forms of feedback, such as numerical ratings or natural language. Neither do we study quality-quantity tradeoffs for learning from pairwise feedback. We leave this exposition to future work.
@@ A.2 Methods that optimize a surrogate reward function
Proximal Policy Optimization [PPO; 24, 57] is a popular RL algorithm that has been recently used to develop InstructGPT [45] and ChatGPT [42]. When applied to fine-tune LMs with RLHF, PPO maximizes the following KL-regularized objective w.r.t. model parameters $\theta$
@@ C.2 Additional results
We now provide additional results for understanding our pairwise annotators.
@@ E.1 Label noise ablations for simulated annotators
Figure 12 sbows the results of the ablation, demonstrating clearly that added label noise provides the majority of the overoptimization effect. In particular, the two options that do not add label noise, $p_{\text{sim}}^{\text{GPT-4}}$ and $p_{\text{sim}}^{\text{eval}}$, keep increasing win-rates with more samples. This result suggests that modeling intra-annotator variability via label noise may be an important component to understanding learning from human preference data.
@@ E.3 Additional analysis of model outputs
Table 7: Average length of the model outputs of baselines and LPF models trained with simulated preferences. We observe that LPF training generally leads to longer outputs.s
@@ Appendix F Additional Analysis on Training and Evaluation Instructions
3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 08 Jan 2024 v4

@@ 4 Validating the AlpacaFarm simulator
Figure 3: The ranking of methods trained and evaluated in AlpacaFarm matches that of methods trained and evaluated in the human-based pipeline. Each point represents one method $M$ (e.g. PPO). The x-axis shows the simulated evaluation (win-rates measured by $p_{\text{sim}}^{\text{eval}}$) on methods trained in simulation $M_{\text{sim}}$. The y-axis shows human evaluation (win-rates measured by $p_{\text{human}}$) on methods trained with human feedback $M_{\text{human}}$. Gray points show models that we did not train, so their $x$ and $y$ values only differ in the evaluation (simulated vs human). Without those points, we have $R^{2}=0.83$ and a Spearman Correlation of $0.94$.
@@ 4.1 Experimental details
As a baseline and starting point for LPF methods, we fine-tuned LLaMA 7B (Touvron23, ) on the 10k SFT split. We take SFT 10k to be the starting point for all LPF methods and collect the simulated preference $p_{\text{sim}}^{\text{ann}}$ and human preference $p_{\text{human}}$ from SFT 10k’s outputs (with temp=1.0) on the 10k instruction PREF split. Then, for each of the six reference LPF methods $M$:
@@
  • We trained some of the models $M$ on human preferences across hyperparameter ranges identified in simulation, evaluating the resulting model $M_{\text{human}}$ against Davinci003 with humans $p_{\text{human}}$.
In addition to the six methods, we also evaluate existing standard instruction following and base models: GPT-4 (gpt-4-0314), ChatGPT (gpt-3.5-turbo-0301), Davinci001 (text-davinci-001), LLaMA 7B (Touvron23, ), and Alpaca 7B (Taori23, ). Alpaca 7B is a LLaMA 7B model finetuned on the concatenation of all data splits, denoted SFT 52k. For these models, we measure both the simulated win-rate and human win-rate using $p_{\text{sim}}^{\text{eval}}$ and $p_{\text{human}}$respectively.
At inference time, for all systems except best-of-$n$, we sample with temperature 0.7 and set the maximum number of tokens to be 300. For best-of-$n$ sampling, we found a higher temperature to be helpful in encouraging output diversity, and so we rerank samples from SFT 10k with temperature 1.0. We provide more thorough experimental details and hyperparameters for all methods in Appendix B.
@@
We collected reference human annotation by showing crowdworkers two potential outputs $y_{0}$ or $y_{1}$ for a given instruction $x$ and asked them to select the index $z\in\{0,1\}$ of their preferred output. Annotators are recruited from Amazon Mechanical Turk using a qualification test of 25 questions. Out of an initial pool of 34 annotators, we selected the 16 whose agreement rate was higher than 70% with the authors’ annotations. We paid the annotators a median hourly rate of \$21, leading to a one-time \$3000 cost of annotating our PREF split and a recurring \$242 cost for evaluating a single model on the 805 evaluation instructions. See Appendix D for additional details including the annotation interface we provided.
@@ 4.2 End-to-end validation of AlpacaFarm
Inspecting these results more closely, we point out the two rank mismatches. The first comparison is SFT10k against SFT52k, where human annotators preferred SFT10k (44.3% vs 40.7%) while the simulator had the opposite preference (36.7% vs 39.2%, Table 2). The other mismatch is ChatGPT against PPO, where human annotators preferred PPO to ChatGPT (55.1% vs 52.9%) unlike the simulator annotators which preferred ChatGPT to PPO (46.8% vs 61.4%). In both cases, these are not major mistakes, as we do not expect SFT52k to be much worse than SFT10k or for a 7B LLaMA model to substantially outperform ChatGPT.
@@ 4.3 Validating the pairwise preferences component
We now show that modeling annotator variability in the simulated annotator ($p_{\text{sim}}^{\text{ann}}$) is necessary to capture important qualitative features of LPF model training. To do so, we compare the behavior of three of the best-performing models trained under $p_{\text{sim}}^{\text{ann}}$ with those trained using the single GPT-4 prompt $p_{\text{sim}}^{\text{GPT-4}}$, which has higher human agreement but little inter- and intra-annotator variability.
Figure 5(c) shows the learning dynamics of these models for pairwise feedback by $p_{\text{human}}$ (left), $p_{\text{sim}}^{\text{ann}}$ (middle), and $p_{\text{sim}}^{\text{GPT-4}}$ (right) as the PPO iteration count and rerank sample counts for best-of-$n$ and expert iteration are increased. For both the human and AlpacaFarm preferences, models that more effectively optimize the surrogate reward (x-axis) improve in win-rate (y-axis) up until a point, where overoptimization of the reward occurs and win-rates decrease. In contrast, simple GPT-4 feedback shows no overoptimization, leading to the false conclusion that the LPF methods can be optimized for much longer than is actually the case. For example, Figure 5(c) right shows best-of-$1024$ to be much better than PPO, which strongly disagrees with the results on human data.
Noted in prior work Gao22 , overoptimization is the result of the proxy reward model $\hat{R}_{\phi}$ being an imperfect estimate of the true reward $R$. We hypothesize that the strong overoptimization seen for human annotators is partly due to the (inter- and intra-) variability of their annotations, which degrades the quality of the reward model. To test this hypothesis, we measured the variance of each of the three annotators by calculating the average error of a held-out annotation to the majority vote over 3 random draws. At 0.26 for $p_{\text{sim}}^{\text{eval}}$ and 0.43 for $p_{\text{sim}}^{\text{ann}}$, AlpacaFarm is close to the human variance of 0.35; in contrast, the GPT-4 annotator has a much lower variance at 0.1. Finally, in Section E.1 we ablate the AlpacaFarm design more finely and find that overoptimization is strongest when using both label noise and 13 annotators, but also seems present when using only label noise. Appendix C contains further analyses of bias and variability of annotators.
@@ 4.4 Validating the evaluation protocol
Finally, we test our evaluation instructions that combines existing open-source evaluation datasets. While we have observed that this dataset is diverse (see Figure 2), it is unclear whether it evaluates performance for any type of real-world human usage. To resolve this, we measure method-level correlations against a set of real user interactions recorded on the Alpaca Demo (Taori23, ). We manually went through the interactions and identified 200 instructions that do not contain any personal identifying information, toxic or unsafe questions, and those that refer to the chatbot directly (e.g. “who are you developed by?”). The terms of use for the demo do not allow us to publicly release this data, but we use this data to evaluate the proposed evaluation set.
@@ 5 Benchmarking reference methods on the AlpacaFarm
Table 2: AlpacaFarm evaluation results on baseline and LHF methods. For methods without a *, the left column shows win-rates when we train and evaluate in simulation, while the right column shows win-rates when we train and evaluate with human feedback. For models with a *, those are not trained by us so the left and right columns respectively show simulated and human evaluation. Win-rates are computed against Davinci003. We did not evaluate Quark and Binary Reward Conditioning for human evaluation because they underperformed in development. ChatGPT and GPT-4 were prompted for responses shorter than 1000 characters to have lengths comparable to other methods.
@@
Davinci001* $24.4\pm 1.5$ $32.5\pm 1.6$
LLaMA 7B* $11.3\pm 1.1$ $6.5\pm 0.9$
We now study the performance of reference methods in AlpacaFarm. Table 2 contains the details of the main evaluation results (presented graphically in Figure 3). In the rest of this section, we discuss our findings from these results, demonstrating that the conclusions we reach using human feedback could have been derived using AlpacaFarm at a substantially lower cost. See appendix B for experimental details.
@@ 5.1 Comparing LPF methods
Among the LPF methods we study, PPO performs the best in both the simulator ($47\%$) and on human feedback data ($55\%$). Notably, with a win-rate of $55\%$, the PPO trained with human feedback was preferred to ChatGPT by our human annotators for single-turn instruction-following. This surprising result is likely because ChatGPT was prompted for responses shorter than 1000 characters to fit in the simulated annotators’ context size, and both crowdworkers and simulators annotators preferred longer responses (see section C.2). For example, in AlpacaEval alpaca_eval (slightly different simulated annotator) we show that the unconstrained ChatGPT achieves $89\%$ win-rate, much higher than the $44\%$ win-rate for the PPO model.
@@
Even though expert iteration trains with best-of-$n$ outputs, it trails best-of-$n$ with a 3–6% win-rate gap in both simulated and human feedback workflows. This result suggests that acquiring the improvement from best-of-$n$ sampling is not straightforward and points to why more complex learning algorithms like PPO might be useful. For Quark, although the rewards of samples improve during training, the resulting models did not outperform the SFT baseline.
@@ 6 Related work
Constitutional AI Bai22 simulates human feedback with AI feedback for model development to improve harmlessness and helpfulness. AlpacaFarm, on the other hand, simulates human feedback with oracle API LLMs so that simulated experiments reflect the outcomes of experiments performed with real human feedback. Due to the difference in goals, the construction and usage of the feedback simulator are different in the two settings. For example, AlpacaFarm’s simulator perturbs LLM preferences with label noise to mimic the noisiness of human annotation, whereas Constitutional AI’s simulator does not inject extra noise.
@@ 7 Limitations and future directions
Validation.
Section 4 validates the use of AlpacaFarm, but there are nevertheless some limitations with the validation setting. First, the instructions we considered (even those from the real world-demo in section 4.4) are relatively simple and single-turn. Second, all models we fine-tuned use a LLaMA 7B as starting point. Finally, human validation is based on feedback from 13 crowdworkers, which may not reflect broader human preferences and seem to have biases such as preferences for longer outputs as highlighted in section C.2. Although we do not expect the previous limitations to significantly affect the usefulness of AlpacaFarm, we encourage users to be vigilant when using any simulator and hope to further validate AlpacaFarm as more datasets and base models become available. More generally, evaluations using pairwise comparison can be an issue even using human annotators as it does not provide a deep understanding of the differences between two models.
Assumption of an oracle LLM.
AlpacaFarm assumes that we can access an oracle LLM, much more powerful than the investigated models, that can be used to approximate human feedback. While this may be true in research settings, this will not be the case when building state-of-the-art models. We believe that investigating the use of the same model for the simulator and for learning is a promising direction for future work.
AlpacaFarm for fine-grained development.
Section 4 shows that AlpacaFarm is effective for model selection and can replicate important behaviors such as over-optimization. We have nevertheless found that suitable hyperparameters for learning algorithms are different for training with simulated feedback compared to human feedback. For example, due to changes in the scale of values of the surrogate reward model, the range of suitable KL regularization coefficients for RLHF is different.
Simulated annotators.
A key component of AlpacaFarm is the simulated annotators. In section 4.3 and appendix C we showed that our annotators have high agreement with human annotators and exhibit similar biases, such as preferences for longer outputs and lists. There are nevertheless some biases that are specific to simulated annotators. For example, we found that simulated annotators prefer the first outputs shown in the prompt and outputs that come from the same model as the simulator. Although we controlled for those biases by randomizing the order of output in the prompt, and using the same training data for all considered methods, there may be other biases that we have not identified. For better automatic annotators and further analysis refer to AlpacaEval alpaca_eval .
@@ A.2 Methods that optimize a surrogate reward function
Proximal Policy Optimization [PPO; 24, 58] is a popular RL algorithm that has been recently used to develop InstructGPT [46] and ChatGPT [43]. When applied to fine-tune LMs with RLHF, PPO maximizes the following KL-regularized objective w.r.t. model parameters $\theta$
@@ C.2 Additional results
We now provide additional results for understanding our pairwise annotators. For more results and code for generating plots see https://github.com/tatsu-lab/alpaca_eval.
@@ E.1 Label noise ablations for simulated annotators
Figure 12 shows the results of the ablation, demonstrating clearly that added label noise provides the majority of the overoptimization effect. In particular, the two options that do not add label noise, $p_{\text{sim}}^{\text{GPT-4}}$ and $p_{\text{sim}}^{\text{eval}}$, keep increasing win-rates with more samples. This result suggests that modeling intra-annotator variability via label noise may be an important component to understand learning from human preference data. The exact ratio of 25% label noise was chosen to be the smallest considered ratio that induced overoptimization out of the ones we considered (0,17%,25%,35%), as shown in fig. 13.
Figure 13: 25% label flipping is the smallest ratio we tested that induced over-optimization. The $y$-axis shows the simulated win-rate. The $x$-axis shows the number $n$ of samples used in best-of-$n$ sampling, which is proportional to the output’s average reward. Each curve shows a reward model trained on samples with different label flipping ratios (0%, 17.5%, 25%, 35%). We see that over-optimization seems to be approximately monotonic with respect to the label-flipping ratio.
@@ E.3 Additional analysis of model outputs
Table 7: Average length of the model outputs of baselines and LPF models trained with simulated preferences. We observe that LPF training generally leads to longer outputs
@@
Table 7: Average length of the model outputs of baselines and LPF models trained with simulated preferences. We observe that LPF training generally leads to longer outputs